OpenStudy (anonymous):

why will noone help me? graph , i graphed top one , i don't know how to do bottom x + y is greater than or equal to –6 2x – 3y < 0

OpenStudy (anonymous):

i have graphed the top line and it looks like this. i don't know how to graph the bottom line

OpenStudy (anonymous):



OpenStudy (queelius):

2x - 3y < 0... Well, there are many ways you can go about this. One way is to express this in a standard line equation, y = mx + b (from 2x - 3y = 0). Another way, and in my this is simpler, is to just pick two different points that satisfy the equation (2x - 3y = 0) and draw a line to connect them. Then, make sure you understand which side of the line satisfies the less than comparison (note: anything actually on the line doesn't satisfy it, since it's less than not less than or equal to).